OpenStudy (anonymous):

lee is taking some friends on a picnic. They'll need to follow a path to get to the picnic spot, a map of the scale is based on a scale of 1:20,000 in cm. If the path is 12 cm on the map how many kilometers is the actual path?

OpenStudy (anonymous):

\[\frac{ 1 }{ 20,000 } = \frac{ 12 }{ x }\]-------------- Cross multiply. \[20,000 \times 12 = 240,000\]So\[1 \times x = 240,000\] 1 times 240,000 equals 240,000. x = 240,000 There are 240,000 kilometers in the actual path.